# ”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder

### v1

@@ 4.2.3. Active Treatment via ERP
Issues with ERP. Overall, all participants were confident about the effectiveness of ERP. However, ERP was less effective in managing mental compulsions (T1, T2). T1 emphasized the limitation of ERP when clients compensated for their behavioral change with mental compulsions, “ They (clients) are doing mental compulsions to sort of cancel out the fact that they didn’t do the physical compulsion, like, okay, I didn’t [check] that doorknob. Okay, I’m making my therapist happy, but what I’m doing in my mind is, I’m going over my mantra 15 times, because that makes me safe. ”
Our findings revealed the thorough process of in-session ERP, from building exposure hierarchy, to exposure design and ad-hoc intensity adjustment, providing insights into designing personalized and flexible in-situ technological support for exposure practices.
4.2.4. Making Long-Term Progress
To maintain long-term progress, professionals highlighted suitable ways to use certain strategies, assigned homework outside therapy sessions, and recommended self-management technologies.
@@ 4.2.4. Making Long-Term Progress
Using Relaxation Techniques Suitably. Participants agreed that relaxation techniques should be taught to clients with high anxiety level, but advised using them only after exposure practices, as clients are expected to sit with their anxiety rather than resolve it immediately (T1, T2, T5, T6). If not used correctly, relaxation techniques can become a new compulsion (T1, T2, T5). As a result, T7 chose not to teach relaxation techniques because they might exaggerate clients’ anxiety if they fail to work, “At one point, [the client] may be really distressed. And they’ve been practicing these breathing techniques that are not going to work for you, then it’s even more distressing, because the only thing you know how to do [with your OCD] is not gonna work.”
Promoting Out-of-Session Success via Homework. Homework is crucial for successful OCD treatment. Participants assigned homework for clients to practice both planned and natural exposures (T1-7) outside of sessions (Section 3.2.3), where clients were asked to record their compulsion duration, anxiety level, and time needed for anxiety to decrease during each practice (T2-T7). Along with exposure practices, clients were instructed to implement learned strategies to promote cognitive change, such as negative thought reframing (T5, T7), ACT-related strategies (e.g., uncertainty mindset) (T1-3, T5-7), which overlap with self-management strategies mentioned in Study 1.
By tracking their compulsive behavior and anxiety level, clients became more aware of their progress and achievements across therapy sessions (e.g., spending less time washing hands). Such positive reflections on progress can enhance confidence and perseverance (T2-4, T6). However, participants also recognized that self-tracking can be difficult to maintain, due to forgetfulness and other daily responsibilities (T3, T6). In addition, exposure practice itself can be overwhelming with multiple triggers present out-of-session, leading to low homework compliance (T3, T7). To address these issues, participants would identify specific barriers faced by individual client, such as fear of exposure (T1, T3), and adjust the homework content accordingly. Although monitoring clients’ progress seemed impossible out-of-session, participants sometimes had clients fill out exposure tracking forms and set up phone reminders to ensure homework completion (T3).
@@
Technology Recommendation for Self-Symptom Management. To ensure clients implementing therapeutic advice outside of therapy, participants also recommended tools and technologies to help clients manage OCD at home (T2-5, T7). Dedicated OCD symptom tracking apps, such as OCDfeat (OCDfeat, 2024), were suggested to easily record compulsions and exposure homework (T2). They also recommended meditation apps, such as Calm (Calm, 2024), for clients to practice mindfulness (T5) and listen to the bilateral stimulation playlist to calm the nervous system (T4) (Kohl et al., 2014). These tools are intended for consistent out-of-session use to monitor symptoms and reinforce skills learned from the therapy.
@@
Our participants also recommended apps for in-situ use in response to OCD episodes upon triggered. To reduce reassurance seeking behavior using specific apps (e.g., web browser), T3 suggested app redirection tool (e.g., ScreenZen (ScreenZen, 2024)) to limit access to those apps on clients’ phones. Interestingly, he also recommended text replacement feature to clients to prevent them from searching specific keywords for obsession-related reassurance, “ I’ll go in and have people create shortcuts where… if they’re going to Google ‘heart attack,’ it actually will try to replace ‘heart [attack]’ with like ‘puppies’ or ‘cute cat videos,’ or something like that.”
However, participants found low engagement in recommended technologies among clients, especially for tracking purpose (T2, T3). While T3 attributed the issue to poor usability of the outdated interface, T2 identified privacy concerns for clients whose OCD symptoms involve taboo thoughts. As she explained “ They (clients) do worry about putting that (taboo thoughts), even writing it down on paper… so they need a lot of reassurance about where that data is being stored and who has access to it.”
@@
When talking about self-administered ERP via mobile apps (e.g., NOCD (NOCD, 2024)), participants recognized its merit in supporting users with low severity (T1-4). However, participants emphasized the importance of psychoeducation in OCD and acquiring adequate symptom management ability via in-person sessions before using such self-help tools (T1, T3, T4, T6, T7).
In general, participants highlighted the need for a tracking tool that preserves privacy. Inspired by participants’ strategies to facilitate long-term success, future self-management technology should guide users to choose strategies wisely, and incorporate progress reflection to maintain high motivation.
@@ 5. Discussion
As an important part of treatment process, self-management of OCD presents significant challenges. The needs for technological support has long been under explored. Through our two studies with people with OCD and the OCD professionals, we sought to triangulate the current gaps and opportunities in OCD symptom self-management. From both studies, we found that people with OCD employed strategies learned from mainstream evidence-based therapies such as ERP and ACT and developed experience-based strategies to resist the urge of compulsions, aligning with professionals’ advice. This confirms the efficacy and generalizability of strategies in OCD therapy.
However, the two studies also highlighted gaps in OCD self-management outside of therapy despite currently-available resources. First, while natural exposure practice is encouraged due to their abundant presence in people’s daily life, these practices were often difficult to complete and tended to be avoided, confirming prior research (Seibell and Hollander, 2014; Eddy et al., 2004). Besides the difficulty of facing the fear, identifying mental compulsions could be challenging and required adequate insights. The absence of physical compulsion does not guarantee the absence of mental compulsions despite going through ERP. Second, some experience-based strategies adopted by people with OCD (Study 1) were identified to be counter-therapeutic and suggested to be prohibited (e.g. reassurance) or used cautiously (e.g., relaxation techniques) by the professionals due to the potential of developing reliance (Study 2). Third, both OCD clients and therapists recognized the importance of symptom and activity tracking, however, traditional tracking tools present limitations and raise privacy concerns due to vulnerability and shame associated with OCD, resulting in the low technology adoption in Study 1 and echoing OCD therapists’ observations in Study 2. Finally, people with OCD faced persisting challenges in OCD self-management despite receiving therapy, highlighting the need for additional support to facilitate smooth social interactions.
The insights from both studies emphasized the needs for self-management technologies beyond formal therapy sessions. By triangulating the insights from people with OCD and the professionals, we identify diverse OCD triggers and properties, compulsive behaviors in relation to triggers, and various effective strategies to inspire the input (e.g., trigger and compulsion recognition) and output (i.e., feedback design) for in-situ self-management technologies for OCD. Furthermore, given the persisting challenges experienced by OCD people, we discuss how technology can address OCD-specific privacy issues and opportunities to support OCD people’s social needs.
@@ 5.1. Enhancing Symptom Recognition & Tracking with User Input
As an important step of insight establishing and symptom self-management, our studies underscored a need for supporting personalized, automatic, ad fine-grained tracking of OCD symptoms and exposure practices.
@@ 5.1.1. Supporting Personalized & Automatic OCD Tracking
From Study 1, we found that OCD triggers are highly individualized, taking diverse forms (e.g., objects, contexts, sounds, thoughts), and that their properties significantly influence trigger intensity. While symptom tracking is essential for improving insights, it is difficult to maintain without therapist supervision (Study 2). Therefore, future technology should incorporate OCD users’ input, allowing specification of triggers to support personalized and automatic trigger detection. Similarly, the system can prompt for the user’s compulsive behaviors based on different types of compulsions (Section 3.2.2) and criteria proposed by OCD therapists (Section 4.2.1). Recent advances in AI have made it possible to adapt pre-trained models to specific domains through few-shot learning (Parnami and Lee, 2022; Zhang et al., 2023). By leveraging just a few examples provided by users, these AI models can be personalized to effectively detect and track OCD symptoms within each user’s unique context. Given the varied form of triggers and compulsions (e.g., direct trigger interaction, consequence mitigation), multi-modal AI models (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024) can be used together with wearable devices (e.g., egocentric camera and microphone (Liu et al., 2024)) to identify and track triggers and compulsions beyond the scope of visual-based methods. Despite the merit of AI-based symptom tracking automation, false positive prediction results can potentially exacerbate OCD symptoms, since people with OCD often have low confidence in their memories (Radomsky et al., 2014). To minimize this risk, prediction models should communicate potential errors and involve users in assessing the decision-making process using explainable AI techniques (Hossain et al., 2023).
@@ 5.1.2. Fine-Grained Symptom Formulation
Besides highly personalized triggers, our findings suggest the importance of considering different properties of triggers to enable fine-grained exposure hierarchy building and symptom tracking. In complement to personalized trigger specification, future tracking technology should prompt the user to reflect on how trigger intensity is affected by different properties of that trigger (Section 3.2.1), measured with SUDS or urge to perform compulsion (Section 4.2.3). As such, an exposure hierarchy with fine-grained and structured trigger properties can be built. During tracking, the system can log fine-grained details of triggers and situate them in the hierarchy as well as identify physical compulsive behaviors, to better support the tracking of exposure practices.
@@
Figure 1a shows a usage scenario that can be achieved with proposed system: the user identified knife as a trigger of his harm OCD, they then are prompted to answer how each property can affect the trigger intensity, for example, a knife can be more triggering when the knife is closer to him, so they assign ‘knife in hand’ with the highest intensity level, and ‘knife on countertop’ with a moderate intensity level. They are also prompted to describe their compulsion—‘avoiding the knife.’ The system will then detect the physical triggers and compulsions in real-time from the camera feed.
For mental compulsions which are difficult to recognize, potential technology can incorporate physiological sensing (e.g., EEG, ECG, eye movement) to detect mental stress (Xia et al., 2019; Al-Jumaily et al., 2021) and mental workload (Fan et al., 2020; Wang et al., 2023), inspired by OCD therapists’ strategies in Study 2. Prior research has shown great potential of using physiological sensing as a therapeutic approach to detect users’ internal states for various mental health conditions. For example, estimating the anxiety level during exposure therapy for anxiety disorder (Gazi et al., 2021; Šalkevicius et al., 2019), and detecting the presence of depression to aid clinical diagnosis(Saeedi et al., 2021; Ay et al., 2019; Thoduparambil et al., 2020). These methods can be adopted to support OCD self-management. However, context information should be incorporated with physiological data since both our work and prior literature (Llorens-Aguilar et al., 2021) indicated the importance of context in recognizing OCD symptoms.
Overall, the proposed technology can potentially improve tracking compliance and precision by formulating fine-grained exposure hierarchy through natural language, and automating OCD symptom tracking using state-of-the-art AI systems. However, we do recognize that our proposed technology could not capture the same level of detail for tracking mental triggers (e.g., worries).
5.2. In-Situ Feedback Design for OCD Self-Management
Consistent with prior literature (Lind et al., 2013; Hezel and Simpson, 2019; Ojalehto et al., 2020), exposure practices can be difficult to follow through outside of therapy sessions, due to the existence of fearful triggers in diverse context. Triggers can appear anywhere anytime and provoke users’ OCD spontaneously, the burden of dealing with unpredicted triggers and resisting compulsions thus all fall onto the users with OCD, potentially resulting in severe consequences. We thus propose the following technologies that provide in-situ intervention during high intensity exposures in people’s everyday life.
5.2.1. In-Situ Compulsion Prevention
Inspired by participants’ strategies in Study 1, and OCD therapists’ technology recommendation, future intervention technology should consider using restrictive interactions to make compulsions more difficult to perform. In the field of human-computer interaction (HCI), researchers have explored the design of such techniques to prevent smartphone overuse (Orzikulova et al., 2024; Lu et al., 2024; Xu et al., 2022), such as delaying a user’s tap for a certain period before it takes effect (Lu et al., 2024). In OCD context, these restrictive interactions can be adapted to help reducing compulsions in digital space, e.g., seeking reassurance online. For compulsive behaviors that happen in physical world, potential technology can explore the design of mixed reality (MR) -based interaction to suppress the urge for compulsions. Researchers has explored the usage of virtual barriers to ensure user safety in MR (Hoang et al., 2022; Sousa et al., 2019). Such design can be adapted for compulsion prevention. Since most overt compulsions involve visual examining, disturbing the user’s vision can effectively interrupt compulsions.
@@ 5.2.1. In-Situ Compulsion Prevention
Figure 1. (a) Fine-Grained Symptom Formulation: an example of the user specifying triggers and trigger properties to build fine-grained exposure hierarchy; (b) AR Compulsion Intervention: an example of using visual augmentation (blurry effect) to disturb the user’s vision.
Figure 1b describes a scenario where the user with contamination OCD washes their hands after contacting triggers (e.g., dirty dishes). Once the system detects excessive hand washing behaviors (longer than 20s), it will blur out or render distracting patterns on the users’ hands and faucet — objects involved in hand washing compulsions. As such, the user’s vision becomes too low to examine their hands, and eventually stopped washing them. Adopting the idea of allowing milder compulsion (Section 4.2.3) when exposure becomes too intense, the invasiveness (e.g., blurry effect intensity) of the visual augmentation should be adjustable. Such intervention can be promising for direct trigger interaction and consequence mitigation that involves active interaction between the user and surrounding physical objects.

### v2

@@ 4.3.2. Compulsion Prevention via Real Life Exposure Practices
What is Missing in Exposure Practice? While all participants were confident about the effectiveness of exposure practices, many acknowledged their limitations in addressing mental compulsions (P2, P5, T1, T2). These covert compulsions, such as self-reassurance or repeated mantras, often persisted even when physical compulsions were reduced. T1 highlighted this challenge, explaining that some clients might compensate for their behavioral change with mental compulsions, “ They (clients) are doing mental compulsions to sort of cancel out the fact that they didn’t do the physical compulsion, like, okay, I didn’t [check] that doorknob. Okay, I’m making my therapist happy, but what I’m doing in my mind is, I’m going over my mantra 15 times, because that makes me safe.” To compensate this limitation, some professionals identified clients’ engagement of mental compulsion alongside their physical compulsive behaviors. For example, T2 described asking clients to rate their mental effort during exposure practices to capture these hidden patterns.
Inspiration:
Inspiration: This highlights the importance of detecting and monitoring mental compulsions in the design of just-in-time intervention technologies to reveal the full picture of a user’s compulsions.
4.3.3. Distraction & Relaxation as Safety Net
When their OCD became so overwhelming that all above strategies failed, OCD participants (P1, P3-P5, P8, P10) turned to self-distraction as a temporary escape—shifting their attention to casual activities such as watching TV (P3), listening to music (P8), calling a friend (P1), or going out for a walk (P5). Similarly, relaxation techniques were used to ease anxiety when breaking the obsession–compulsion cycle felt impossible (P1, P3, P6, P7, P10). Such strategies included breathing techniques, muscle relaxation, and grounding exercises.
@@ 4.3.3. Distraction & Relaxation as Safety Net
While professionals supported using these methods to manage intense anxiety (T1, T2), they cautioned that such strategies must be used carefully. They noted the importance of helping clients learn to tolerate anxiety rather than eliminate it (T1, T2, T5, T6). Without careful guidance, overreliance on distraction or relaxation can itself become a form of compulsion (T1, T2, T5). As a result, T7 chose not to teach relaxation techniques at all because they might backfire if they don’t work when clients need them most, “At one point, [the client] may be really distressed. And they’ve been practicing these breathing techniques that are not going to work for you, then it’s even more distressing, because the only thing you know how to do [with your OCD] is not gonna work.”
Inspiration:
Inspiration: Relaxation strategies can be beneficial to people with OCD only when used correctly. It is important for just-in-time technology to involve proper guidance to support and prevent relaxation at the right timing.
4.3.4. External Mental Health Support
Psychotherapy is a key resource for learning OCD coping strategies (O’Neill and Feusner, 2015). While most participants reported positive experiences, some felt lost and discouraged in therapy not specialized for OCD (e.g., generic talking therapy for P6). Nine out of 10 participants also experienced challenges accessing OCD therapy due to long wait time (P1, P5–P7, P9, P10), high costs (P1, P3–P5, P7–P9), and stigma around OCD (P6, P7), underscoring the need for supplementary forms of support.
Beyond formal therapy treatment, most participants (P1–8, P10) received support from family and friends to help manage their symptoms. Some had close contacts who understood their condition and followed therapeutic guidance—helping monitor symptoms (P4) and encouraging them to challenge their OCD independently (P1–3, P7).
Others, however, had family members or friends who misunderstood OCD and unintentionally reinforced symptoms by providing accommodations contrary to therapists’ suggestions. For example, P6’s mother washed her hands every time after using elevators to accommodate P6’s contamination-related fears. As several professionals emphasized (T1–5), psychoeducation on how therapy works and the expected outcome of OCD therapy for family members is critical, so they can respond to their loved ones in ways that support, not hinder, treatment progress (T2–4).
Inspiration:
Inspiration: Psychoeducation could be integrated into OCD self-management technologies to improve awareness and collaboration between people with OCD and their support system.
4.4. Opportunities and Gaps in OCD Self-Management Technologies
In this section, we explore participants’ experiences and attitudes toward existing technologies, and identified key challenges in OCD self-management despite the availability of technological support. Based on this, we discuss opportunities to better support OCD self-management.
4.4.1. Technology Recommendation for OCD Self-Management
@@
Surprisingly, none of our OCD participants reported using technology specifically designed for mental health support. In contrast, OCD professionals recommended a range of tools from symptom tracking to compulsion prevention to help clients manage their OCD outside of therpy (T2-5, T7). Dedicated OCD symptom tracking apps, such as OCDfeat (OCDfeat, 2024), were suggested to easily record compulsions and exposure practices (T2). They also recommended meditation apps, such as Calm (Calm, 2024), for clients to practice mindfulness (T5) and listen to the bilateral stimulation playlist to calm the nervous system (T4) (Kohl et al., 2014). These tools are intended for consistent out-of-session use to monitor symptoms and reinforce skills learned from the therapy.
@@
Professionals also recommended apps for in-situ use in response to OCD episodes upon triggered. To reduce reassurance seeking behavior using specific apps (e.g., web browser), T3 suggested app redirection tool (e.g., ScreenZen (ScreenZen, 2024)) to limit access to those apps on clients’ phones. Interestingly, he also recommended text replacement feature to clients to prevent them from searching specific keywords for obsession-related reassurance, “ I’ll have people create shortcuts where… if they’re going to Google ‘heart attack,’ it actually will try to replace ‘heart [attack]’ with like ‘puppies’ or ‘cute cat videos,’ or something like that.”
@@
When talking about self-administered ERP via mobile apps (e.g., NOCD (NOCD, 2024)), professionals recognized its merit in supporting users with low severity (T1-4). However, they emphasized the importance of psychoeducation in OCD and acquiring adequate symptom management ability via in-person sessions before using such self-help tools (T1, T3, T4, T6, T7).
Inspiration:
Inspiration: To prevent overreliance on technology becoming a new compulsion, just-in-time intervention should adopt a wind-down strategy to gradually help users be independent on symptom management.
4.4.2. Inconvenience & Privacy Concerns in Symptom Tracking
Daily symptom tracking helps people with OCD build self-awareness around their behaviors and thoughts (P1, P4, P5, P8). It also allows them to observe progress and recognize achievements (e.g., spending less time doing compulsions) over time, which can foster a sense of accomplishment, increase confidence, and strengthen motivation to continue practicing coping strategies (T2–T4, T6). Participants typically logged daily triggers (P4, P8), their responses to those triggers (P1, P4), and distress level during exposure practices (P1, P5) either physically or digitally (e.g., Google Doc).
However, maintaining a consistent tracking habit was challenging (P1, P2, P7, P8, T3, T6). OCD participants found traditional journaling methods superficial, as it failed to provide meaningful insights into participants’ symptom patterns or suggest optimal management strategies (P1, P8). Professionals, on the other hand, attributed low engagement to forgetfulness due to other daily responsibilities (T3, T6). To address this, professionals sometimes had clients fill out symptom tracking forms and set up phone reminders to ensure compliance (T3).
Privacy concerns also hindered the use of symptom tracking tools (P7, T2). Participants were especially cautious about recording sensitive thoughts or exposure practices without clear information on how their data would be stored and protected. P7, for example, expressed worry about the emotional risk of recording vulnerable moments: “I would be most concerned about privacy issues, because it’s very vulnerable for people… to do all of that exposure [practices] in a place where their data is [at stake]. So if sessions are recorded… or their speech or their notes are recorded, making sure all of that is well protected… ” Similarly, T2 noted that clients whose OCD symptoms involved taboo thoughts often avoided logging due to fears about data privacy and access, as she explained, “ They (clients) do worry about putting that (taboo thoughts), even writing it down on paper… so they need a lot of reassurance about where that data is being stored and who has access to it.”
Inspiration:
Inspiration: Symptom tracking is a crucial first step for enabling effective just-in-time interventions. These findings highlight the need for a simple, transparent tool that can automate tracking while respecting and protecting users’ privacy.
4.4.3. OCD-Induced Social Barriers
In addition to challenges with self-management, participants with OCD also described how OCD created difficulties in their social lives. Due to the heterogeneity of participants’ OCD symptoms, understanding the nuanced experience of OCD could be difficult, even with OCD professionals (P6, P10). P10 had to downplay his symptoms to gain empathy from his therapist. Outside of therapy, participants often felt their compulsions interrupted social interactions (P2, P7, P8, P9). For instance, P9 found it embarrassing to go back to check the parking brake of his car during group activities. Moreover, exposure practices in real life can also bring social barriers since practicing in public spaces or when other people are around could be socially awkward (P1, P7).
Inspiration:
Inspiration: More support is needed across social, clinical, and technological contexts to foster greater awareness of OCD and ensure that people with OCD have access to supportive environments and resources.
@@ 5. Discussion
As an important aspect of quality of life, self-management of OCD presents significant challenges and the needs for technological support has long been underexplored. Our research fill this gap by thoroughly investigating the perspectives of both people with OCD and OCD professionals. Our findings centered on the lived experiences of people with OCD and supplemented and contrasted with the professionals’ views on treatment strategies and technology recommendations. As a result, we identified the gaps and opportunities in OCD symptom self-management.
First, while natural exposure practice was encouraged due to their abundant presence in daily life, these practices were often difficult to complete and tended to be avoided (Seibell and Hollander, 2014; Eddy et al., 2004). Besides the difficulty of facing the fear, identifying mental compulsions could be challenging and required adequate insights. The absence of physical compulsion does not guarantee the absence of mental compulsions despite going through ERP. Second, some experience-based strategies adopted by people with OCD (e.g., relaxation techniques) were suggested to be used cautiously by the professionals due to the potential risk of developing reliance. Third, both OCD participants and therapists recognized the importance of symptom and activity tracking, however, traditional tracking tools present limitations and raise privacy concerns due to vulnerability and shame associated with OCD, resulting in low technology adoption. Finally, people with OCD faced persisting challenges in OCD self-management especially in social scenarios despite receiving therapy, highlighting the need for additional support to facilitate smooth social interactions.
These gaps emphasized the needs for self-management technologies beyond formal therapy sessions. Based on the diverse OCD triggers, compulsive behaviors, and symptom management strategies identified in our research, we discuss the input (e.g., trigger and compulsion recognition) and output (i.e., feedback design) of future just-in-time self-management technologies for OCD. Furthermore, given OCD people’s challenges, we discuss how technology can address OCD-specific privacy issues and opportunities to support OCD people’s social needs.
@@ 5.1. Enhancing Symptom Recognition & Tracking with User Input
As an important step of establishing symptom self-awareness for symptom self-management, our study underscored a need for supporting personalized, automatic, and fine-grained tracking of OCD symptoms and exposure practices.
@@ 5.1.1. Supporting Personalized & Automatic OCD Tracking
We found that OCD triggers are highly individualized, taking diverse forms (e.g., objects, contexts, sounds, thoughts), and that their properties significantly influence trigger intensity. While symptom tracking is essential for improving symptom awareness and motivation, it is difficult to maintain without therapist supervision. Therefore, future technology should incorporate OCD users’ input, allowing specification of triggers to support personalized and automatic trigger detection. Similarly, the system can prompt for the user’s compulsive behaviors based on different types of compulsions and criteria proposed by OCD therapists (Section 4.2). Recent advances in AI have made it possible to adapt pre-trained models to specific domains through few-shot learning (Parnami and Lee, 2022; Zhang et al., 2023). By leveraging just a few examples provided by users, these AI models can be personalized to effectively detect and track OCD symptoms within each user’s unique context. Given the varied form of triggers and compulsions (e.g., direct trigger interaction, consequence mitigation), multi-modal AI models (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024) can be used together with wearable devices (e.g., egocentric camera and microphone (Liu et al., 2024)) to identify and track triggers and compulsions beyond the scope of vision-based methods. Despite the merit of AI-based symptom tracking automation, false positive prediction results can potentially exacerbate OCD symptoms, since people with OCD often have low confidence in their memories (Radomsky et al., 2014). To minimize this risk, prediction models should communicate potential errors and involve users in assessing the decision-making process using explainable AI techniques (Hossain et al., 2023).
@@ 5.1.2. Fine-Grained Symptom Formulation
Besides highly personalized triggers, our findings suggest the importance of considering different properties of triggers to enable fine-grained exposure hierarchy building and symptom tracking. In complement to personalized trigger specification, future tracking technology should prompt the user to reflect on how trigger intensity is affected by different properties of that trigger, measured with SUDS or urge to perform compulsion (Section 4.1). As such, an exposure hierarchy with fine-grained and structured trigger properties can be built. During tracking, the system can log fine-grained details of triggers and situate them in the hierarchy as well as identify physical compulsive behaviors, to better support the tracking of exposure practices.
@@
Figure 1 shows a usage scenario that can be achieved with proposed system: the user identified “trash can” as a trigger of their contamination OCD, they then are prompted to answer how each property can affect the trigger intensity, for example, a trash can can be more triggering when the trash can is overflowing, so they assign “full of trash” with the highest intensity level, and “no trash” with a moderate intensity level. They are also prompted to describe their compulsion—“avoiding using the trash can.” The system will then detect the physical triggers and compulsions in real-time from the camera feed.
Mental compulsions are often difficult to recognize but critical to address. To support their detection, future technologies could incorporate physiological sensing, such as EEG and ECG, to assess mental workload (Fan et al., 2020; Wang et al., 2023), inspired by strategies used by professionals (T2). Prior research has also shown promising results in detecting mental rumination using self-report measures combined with wearable sensor data, like EMG (Nalborczyk et al., 2017; Arnold and Young, 2025). By capturing a user’s full compulsion pattern and integrating contextual information (e.g., triggers, environment), the system can generate more tailored and effective intervention strategies.
Despite the potential, prompting users to report symptoms itself can be triggering. Our suggested input methods may require a certain level of symptom awareness and readiness for symptom management, which should be assessed before prompting users’ symptoms. Addressing this challenge is essential for creating safe and effective self-management tools. Future research should explore symptom reporting methods that minimize discomfort while ensuring honesty.
@@
Figure 1. Fine-Grained Symptom Formulation: an example of the user specifying triggers and trigger properties to build fine-grained exposure hierarchy.
[…]
The figure mainly contains three part left to right. On the left, it shows a conversation interface between a user with OCD and an AI agent. The agent asks the user ‘What is your OCD trigger?’ the user answers ‘Trash cans.’ Then the agent says ‘Please tell me how following properties of a knife affect its trigger intensity. Let’s start with Size.’ The user’s speech bubble shows ‘…’ In the middle, there are three sliders labeled with different properties. All three sliders are of same length with the left end representing ‘Lowest SUDS’, and right end representing ‘Highest SUDS’. There are two knobs on the ‘size’ slider saying ”wastebasket at home” and ”big garbage can outside” located at the middle-left and the right end of the slider, respectively. There are two knobs on the ‘Trash inside?’ slider saying ”no trash” (located near the left end of the slider) and ”full of trash” (located on the right end). There are two knobs on the ‘With a lid?’ slider saying ”no lid” (located near the left end of the slider) and ”lid with a handle” (located near the right end). On the right of the figure, it shows an illustration of a silver trash can with a recycling symbol on it.
5.2. Just-in-Time Feedback Design for OCD Self-Management
Consistent with prior literature (Lind et al., 2013; Hezel and Simpson, 2019; Ojalehto et al., 2020), exposure practices can be difficult to follow through outside of therapy sessions, due to the existence of fearful triggers in diverse context. Triggers can appear anywhere anytime and provoke users’ OCD spontaneously. As such, the burden of dealing with unpredictable triggers and resisting compulsions all fall onto the users with OCD. This can potentially result in severe consequences. We thus propose the following technologies that provide just-in-time intervention during high intensity exposures in people’s everyday life.
5.2.1. Augmented Reality-Based Compulsion Prevention Support
Inspired by OCD participants’ strategies and OCD therapists’ technology recommendation, future intervention technology should consider using restrictive interactions to make compulsions more difficult to perform. Inspired by prior work on smartphone overuse prevention (Orzikulova et al., 2024; Lu et al., 2024; Xu et al., 2022), future intervention technology can adopt similar restrictive interactions to help reducing compulsions in digital space, e.g., seeking reassurance online. For compulsive behaviors that happen in physical world, potential technology can explore the design of augmented reality (AR) -based interaction to suppress the urge for compulsions. Researchers has explored the usage of visual barriers in AR and VR (virtual reality) systems to ensure user safety (Hoang et al., 2022; Sousa et al., 2019). Such design can be adapted for compulsion prevention. Since most overt compulsions involve visual examining (e.g., repeated checking), adequately disturbing the user’s vision by placing visual barriers on objects being examined may effectively interrupt physical compulsions. For instance, rendering visual blockage on stovetop when the user repeatedly checking burner knobs.